what is the velocity of a rock which was thrown down from a bridge if it has a velocity of 35 m/s after 3 s?


1
Expert's answer
2021-10-10T16:26:00-0400

We can find the initial velocity of the rock from the kinematic equation (we choose the downward as a positive direction):


"v=v_0+gt,""v_0=v-gt=35\ \dfrac{m}{s}-9.8\ \dfrac{m}{s^2}\cdot3\ s=5.6\ \dfrac{m}{s}."